• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name folts

    Log entry time 08:21:58 on May 15, 2008

    Entry number 238014

    keyword=Hall Status

    Target chamber id closed up and pumped down to 5x10-5

    Right spectrometer power supplies are turned off and locked out with admin locks

    Right spectrometer vacuum pumps are off

    He will be removed from the right spectrometer magnets today